Here are the summary statistics from the two random samples.

sample size sample mean sample standard deviation
Females 50 98.4060 0.6557
Males 50 98.1100 0.6710

Test statistic: z∗=(x¯1−x¯2)s21n1+s22n2√

Compute the test statistic (round off your final answer to two decimal places) and input your answer below.
Answer #1
female male
x1    = 98.4060 x2      = 98.1100
s1    = 0.6557 s2      = 0.6710
n1    = 50 n2     = 50
standard error se=√(S21/n1+S22/n2)=sqrt(0.6557^2/50+0.6710^2/50) = 0.1327

therefore test statistic (x1-x2)/standard error =(98.4060-98.1100)/0.1327 = 2.23